OpenStudy (anonymous):

If ƒ(x)=−10x−6ƒ(x)=−10x−6, then ƒ^−1(x)=

OpenStudy (astrophysics):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (lynfran):

first rewrite the function as an equation y=-10x-6 solve for x y+6=-10x -(y+6)/10=x then write f^-1(y) f^-1(y)=-(y+6)/10 or f^-1(x)=-(x+6)/10 then check.. f(f^-1(x))=-10(f^-1(x))-6 can u finish from here...

OpenStudy (anonymous):

sorry i just realized i typed it wrong

OpenStudy (anonymous):

its if ƒ(x)=−10x−6, then ƒ^−1(x)=

OpenStudy (lynfran):

i know ... thats the same function i did above...

OpenStudy (anonymous):

okay so do i solve that?

OpenStudy (astrophysics):

All the steps are given, please try it out.

OpenStudy (anonymous):

okay thank you